OpenStudy (anonymous):

2691% Convert this percent into a Decimal. A. .002691 B. .2691 C. 2.691 D. 26.91

OpenStudy (anonymous):

can some body help me please?

OpenStudy (anonymous):

26.91

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Look : 2691%=2691*1/100=2691/100

OpenStudy (anonymous):

thanks you so much(:

OpenStudy (anonymous):

welcome :)

OpenStudy (cggurumanjunath):

d